A rectangular floor of a community hall has a width of 1800 cm and a length of 2400 cm. The floor will be covered with wooden tiles. Each tile is a square of side 30 cm. The cost of tiling is RM18 per tile. Calculate the total cost in tiling the floor of the hall.


1
Expert's answer
2021-09-26T17:56:48-0400

Area of the rectangular floor "=1800 \times 2400=4320000\ cm^2"

Area of square tile "=(30)^2=900\ cm^2"

So, total number of tiles "=\frac{4320000}{900}=4800"

The cost of tiling is Rs.18 per tile.

So, overall cost "=4800\times 18=Rs. 86400"